# Unlocking the Capabilities of Vision-Language Models for Generalizable and Explainable Deepfake Detection

### v1

@@ 4.1 Experimental Settings
Datasets. The FaceForensics++ (FF++) (Rossler et al., 2019) dataset includes 1,000 real videos and 5,000 forgery videos across five deepfake categories, which is one of the most widely-used datasets for deepfake detection. DFD (Dufour & Gully, 2019), CDF1 , CDF2 (Li et al., 2020b), DFDCP (Dolhansky, 2019), and DFDC (Dolhansky et al., 2020) are commonly used datasets for evaluating generalization performance in deepfake detection. The images are all cropped using Dlib and RetinaFace. To be consistent with the previous deepfake detection approaches, we trained only on real data from the FF++ dataset.
Evaluation Metrics. Following existing approaches (Shiohara & Yamasaki, 2022; Nguyen et al., 2024), we used the video-level Area Under the Receiver Operating Characteristic Curve (AUC) and Average Precision (AP) as our evaluation metric. Additionally, we assess the LLM’s performance by evaluating the binary classification (Yes or No) of authenticity in its textual output, allowing us to calculate a corresponding video-level AUC for the LLM’s responses.
Compared Methods. We evaluated our approach against several state-of-the-art deepfake detection algorithms (Rossler et al., 2019; Li et al., 2020a; Zhao et al., 2021a; Liu et al., 2021; Zhao et al., 2021b; Cao et al., 2022; Shiohara & Yamasaki, 2022; Wang et al., 2023a, b; Dong et al., 2023; Yan et al., 2023; Qian et al., 2020; Nguyen et al., 2024) and LVLM-Based Methods (Khan & Dang-Nguyen, 2024; Su et al., 2023; Liu et al., 2024b; Wang et al., 2024a).
Implementation Details. Our approach leverages the PandaGPT architecture, which incorporates the ImageBind-Huge model as the image and text encoder. We extract features from the 16th, 24th, and 32nd layers of the encoder to compute consistency maps with the text features, which are then passed to the Vicuna-7B model for inference. For multi-turn dialogue capability, we alternate training between the deepfake dataset and the PandaGPT dataset. All the images are cropped to 224 $\times$ 224. Training is conducted on two Nvidia RTX 4090 GPUs over 50 epochs, using the Adam optimizer with a learning rate of 1e-4 and a weight decay of 1e-5.
@@ 4.2 Comparison with SOTA Detection Methods
Intra-Dataset Evaluation. Following the intra-dataset protocol outlined in (Yan et al., 2024; Nguyen et al., 2024), we compare our approach with existing state-of-the-art deepfake detection methods based on the outputs of KFD. As shown in Table 1, our method achieves competitive results, reaching a detection AUC of 99.53% on the FF++ dataset.
Cross-Dataset Evaluation. Following previous works (Li et al., 2020a; Shiohara & Yamasaki, 2022; Nguyen et al., 2024), we further perform a cross-dataset evaluation. The models are trained on real data from the FF++ dataset, and the detection performance is assessed on CDF1, CDF2, DFD, DFDCP, and DFDC datasets. We report video-level AUC and AP scores across these datasets. The results are summarized in Table 1. Our model surpasses the best existing methods, achieving a performance gain in detection AUC of 2.12%, and 4.18% on the DFDC, and CDF1 datasets, respectively. Overall, our approach provides a mean improvement of 1.34% in AUC, highlighting its generalization ability across diverse datasets.
@@
Figure 6: Forgery localization results and LLM responses on various forgery types from the FF++ dataset. For each example, the top-left shows the original image, the top-right displays the forgery segmentation map, and the text below provides the LLM’s textual detection result.
Methods FF++ CDF2 DFD CDF1 DFDCP Avg P-Tuning 91.32 79.97 88.74 81.45 69.26 82.15 PandaGPT 63.42 62.53 64.56 55.01 46.06 58.32 Qwen2-VL 50.50 48.88 42.38 46.46 51.71 47.99 FAK-Owl 67.54 73.51 68.61 62.13 65.96 67.55 Ours-LLM 97.11 89.90 92.26 95.97 86.70 92.39 Ours-VLM 97.11 89.90 92.26 95.97 86.70 92.39
Table 2: Comparison of our scheme with existing LVLM-based methods. All AUC values are calculated based on textual outputs generated by the LVLM. P-Tuning (Khan & Dang-Nguyen, 2024) and PandaGPT (Su et al., 2023) are fine-tuned on the FF++ dataset, FAK-Owl (Liu et al., 2024b) is trained based on original image-text pairs, and Qwen2-VL (Wang et al., 2024a) is evaluated with its original pre-trained weights. The best and second-best results are highlighted in bold and underlined, respectively.
GradCAM Visualization. In this section, we use GradCAM (Selvaraju et al., 2017) to visualize attention during deepfake detection. As shown in Figure 5, we compare our method with SBI (Shiohara & Yamasaki, 2022) across the CDF2, DFD, and DFDC datasets. While SBI can detect forged regions, it may misidentify these areas and tends to highlight irrelevant regions when encountering unknown forgeries. In contrast, our scheme, guided by external knowledge, effectively identifies regions that are inconsistent with predefined textual descriptions.
@@ 4.3 Comparison with LVLM-based Methods
Detection Performance of LVLM. We further benchmark our framework against leading LVLM-based methods (Zhou et al., 2022; Su et al., 2023; Wang et al., 2024a). In these evaluations, both the image and the query are provided as inputs to the LVLM, with the model tasked to determine the authenticity of the image (real or fake). For PandaGPT, Qwen2-VL, and our scheme, we utilize the LLM’s output (“Yes” or “No”) to classify authenticity, and subsequently compute the AUC values for evaluation. As shown in Table 2, compared to existing LVLM-based approaches, our model shows superior results, achieving a performance improvement of 5% , 9%, 3%, 14% ,and 10% on the FF++, CDF2, DFD, CDF1, and DFDCP datasets, respectively.
Dialogues Visualization. Unlike prior detection methods, our approach not only supports deepfake detection but also facilitates multi-turn dialogues, enabling users to further inquire about the image content. Figure 6 presents some example dialogues on the FF++ dataset. It shows that our proposed scheme accurately identifies forged regions within the images, while the LLM provides precise and contextually relevant judgments. Additional multi-turn dialogue examples are provided in the supplementary material.
@@ 4.4 Analysis
Training Number Test set AUC FF++ CDF1 DFD DFDC Avg 50 92.68 83.18 74.46 61.82 78.04 100 94.38 91.12 78.99 63.88 82.09 200 94.02 91.98 80.47 62.09 82.14 500 95.15 91.82 81.65 65.31 83.48
@@
Number of Training Images. Obtaining large-scale face images in real-world scenarios is often infeasible. Hence, we assess our model’s performance with varying numbers of training images. Specifically, we randomly sample 50, 100, 200, and 500 real images from FF++ to create corresponding fake image-text pairs for training. As shown in Table 3, our approach achieves state-of-the-art performance using only 500 training images. Although reducing the training set size leads to a slight degradation in accuracy, our method still maintains competitive results with as few as 100 training images. This underscores the robustness of our framework, particularly in scenarios where data are scarce.
Effect of Prompt Tuning. The prompt tuning process is designed to convert forgery detection knowledge into the input of LLM to facilitate accurate detection. This process involves the Forgery Prompt Learner (FPL), the LLM, and the LoRA strategy. To evaluate the effectiveness of this component, we conducted ablation experiments on four subsets of FF++ as well as CDF2 and DFDCP datasets. For each configuration, the AUC was calculated based on the LLM’s output in determining real versus fake images. As shown in Table 4, models equipped with the Forgery Prompt Learner demonstrate higher AUC values, indicating improved effectiveness for deepfake detection tasks. Furthermore, the integration of LoRA further enhances performance, achieving superior results across multiple datasets compared to configurations without LoRA.
Training Set Test Set AUC (%) Database #Real FF++ CDF2 DFDCP DFDC Avg FF++ 720 99.54 94.32 92.39 77.03 90.82 CDF2 622 97.95 95.79 86.36 70.27 87.59 DFDCP 737 99.15 85.17 93.74 69.64 86.93
@@
LLM Architecture Test set AUC FF++ CDF1 DFD DFDC Avg Llama-3.2-1B 96.49 95.64 82.74 65.66 85.13 Llama-3.2-3B 97.08 95.30 82.80 66.12 85.32 Vicuna-7B 97.11 95.97 84.26 67.65 86.25
@@
Generalizability to Training Datasets. The generalizability of deepfake detection is closely tied to the diversity of the training data used. To verify the effectiveness of our approach across different datasets, we trained the model on various training sets and conducted cross-dataset evaluation on FF++, CDF2, DFDCP, and DFDC datasets. We calculated detection AUC and AP values based on the LLM’s responses, as shown in Table 5. The results indicate that our method exhibits strong robustness across diverse datasets and demonstrates its ability to generalize across different types of forgeries using varied real data.
Effects of different LLM architectures. The deepfake detection performance is closely related to the specific LLM architecture used, as different LLMs exhibit unique characteristics. To examine detection performance across various LLM architectures, we evaluated three models: Llama-3.2-1B, Llama-3.2-3B, and Vicuna-7B. The evaluation is conducted on the FF++, CDF1, DFD, and DFDC datasets. As shown in Table 6, larger-scale architectures consistently achieved better detection performance, indicating that models with greater parameter counts are better suited to capturing subtle forgery artifacts.
@@ 5 Conclusion
In this work, we introduced a novel deepfake detection framework that leverages Large Vision-Language Models (LVLMs) to enhance generalization and explainability. By integrating a Forgery Vision-Text Alignment Module, we effectively align image features with textual descriptions of pristine and deepfake images to facilitate the generation of forgery consistency maps. Additionally, we incorporated a Forgery Prompt Learner capable of transforming fine-grained forgery features into inputs for the LLM, enabling accurate forgery detection responses. Extensive evaluations across multiple benchmarks, including FF++, CDF1, CDF2, DFD, DFDCP, and DFDC, demonstrate that our scheme outperforms existing methods in generalization performance. These findings underscore the potential of LVLM-based approaches in advancing deepfake detection methodologies.
Appendix A You can have an appendix here.
You can have as much text here as you want. The main body must be at most $8$ pages long. For the final version, one more page can be added. If you want, you can use an appendix like this one.
The \$\mathtt{\backslash onecolumn}$ command above can be kept in place if you prefer a one-column appendix, or can be removed if you prefer a two-column appendix. Apart from this possible change, the style (font size, spacing, margins, page numbering, etc.) should be kept the same as the main body.

### v2

@@ 4.1 Experimental Settings
Datasets. The FaceForensics++ (FF++) (Rossler et al., 2019) dataset includes 1,000 real videos and 5,000 forgery videos across five deepfake categories, which is one of the most widely-used datasets for deepfake detection. DFD (Dufour & Gully, 2019), CDF1, CDF2 (Li et al., 2020b), DFDCP (Dolhansky, 2019), DFDC (Dolhansky et al., 2020), and DF40 (Yan et al., ) are commonly used datasets for evaluating generalization performance in deepfake detection. The images are all cropped using Dlib and RetinaFace. We train only on real data from the FF++ dataset.
Evaluation Metrics. Following existing approaches (Shiohara & Yamasaki, 2022; Nguyen et al., 2024), we use the video-level Area Under the Receiver Operating Characteristic Curve (AUC) and Average Precision (AP) as our evaluation metric. Additionally, we assess the LLM’s performance by evaluating the binary classification (Yes or No) of authenticity in its textual output, allowing us to calculate a corresponding video-level AUC for the LLM’s responses.
Compared Methods. We evaluate our approach against several state-of-the-art deepfake detection algorithms (Rossler et al., 2019; Li et al., 2020a; Qian et al., 2020; Zhao et al., 2021a; Liu et al., 2021; Zhao et al., 2021b; Cao et al., 2022; Shiohara & Yamasaki, 2022; Wang et al., 2023a, b; Dong et al., 2023; Yan et al., 2023; Xu et al., 2023; Yan et al., 2024; Nguyen et al., 2024; Cheng et al., 2024; Lin et al., 2025; Luo et al., 2024; Ba et al., 2024; Fu et al., 2025) and LVLM-based Methods (Khan & Dang-Nguyen, 2024; Su et al., 2023; Liu et al., 2024b; Wang et al., 2024a)
Implementation Details. Our approach leverages the PandaGPT architecture, which incorporates the ImageBind-Huge model as the image and text encoder. We extract features from the 16th, 24th, and 32nd layers of the encoder to compute consistency maps with the text features, which are then passed to the Vicuna-7B model for inference. For multi-turn dialogue capability, we alternate training between the deepfake dataset and the PandaGPT dataset. The forgery region number $n$ is set as 3. All the images are cropped to 224 $\times$ 224. Training is conducted on two Nvidia RTX 4090 GPUs over 50 epochs, using the Adam optimizer with a learning rate of 1e-4 and a weight decay of 1e-5. The loss parameter $\lambda$ is set as 1.
@@ 4.2 Comparison with SOTA Detection Methods
Intra-Dataset Evaluation. Following the intra-dataset protocol outlined in previous works (Yan et al., 2024; Nguyen et al., 2024), we compare our approach with existing state-of-the-art deepfake detection methods based on the outputs of KFD. As shown in Table 1, our method achieves competitive results, reaching a detection AUC of 99.53% on the FF++ dataset.
Cross-Dataset Evaluation. Following previous works (Li et al., 2020a; Shiohara & Yamasaki, 2022; Nguyen et al., 2024), we further perform a cross-dataset evaluation. The models are trained on real data from the FF++ dataset, and the detection performance is evaluated on CDF1, CDF2, DFD, DFDCP, and DFDC datasets. We report video-level AUC and AP scores across these datasets in Table 1. As summarized in Table 1, our method achieves superior video-level AUC and AP scores across these diverse datasets, outperforming state-of-the-art baselines.
DF40-FS UniFace InSwapper FSGAN FaceDancer e4s SBI 89.02 88.52 89.62 78.18 86.36 CADDM 86.86 78.65 88.86 76.54 87.92 Ours 90.61 87.64 93.75 82.97 94.68 DF40-FR PIRender HyperReenact FOMM FS_vid2vid LIA SBI 81.81 65.31 88.05 83.72 89.22 CADDM 77.37 69.26 84.77 72.86 69.67 Ours 88.29 81.55 93.34 71.56 99.99 DF40-EFS StyleGAN2 StyleGAN-XL DDIM DiT-XL/2 PixArt-$\alpha$ SBI 97.91 23.26 99.56 79.04 98.78 CADDM 100.00 98.69 98.10 79.90 99.74 Ours 100.00 100.00 99.93 94.18 100.00
Table 2: Cross-manipulation evaluation on DF40 dataset. All methods are evaluated on various types of manipulated subsets of DF40. The best-performing values are highlighted in bold.
@@
Cross-Manipulation Evaluation of KFD. To evaluate the robustness of our algorithm against unknown forgery techniques, we assess its detection performance on data generated by unseen manipulation methods within the DF40 dataset. DF40 comprises synthetic deepfake samples generated from real images in the FF++ dataset, encompassing a diverse range of manipulation methods. Our models are trained exclusively on real images from FF++ and subsequently tested on unseen manipulation types within DF40. As presented in Table 2, our approach consistently outperforms existing methods across multiple forgery types, achieving state-of-the-art detection results. Notably, our method demonstrates robust generalization capabilities when confronted with advanced deepfake generation techniques, including FSGAN, E4S, LIA, StyleGAN, DDIM, and PixArt-$\alpha$.
GradCAM Visualization. To further interpret the decision process of our model, we use GradCAM (Selvaraju et al., 2017) to visualize attention when encountering unknown datasets. As shown in Figure 4, we compare our method with SBI (Shiohara & Yamasaki, 2022) across the CDF2, DFD, and DFDC datasets. While SBI can detect forged regions, it may misidentify these areas and tends to highlight irrelevant regions when encountering unknown forgeries. In contrast, our scheme, guided by external knowledge, effectively identifies forgery regions.
@@ 4.3 Comparison with LVLM-based Methods
Methods FF++ CDF2 DFD CDF1 DFDCP Avg Clipping 91.32 79.97 88.74 81.45 69.26 82.15 RepDFD - 89.94 - - 95.03 92.49 UDD - 93.13 - - 88.11 90.62 Ours-KFD 99.53 94.71 99.64 97.62 91.81 96.66 PandaGPT 63.42 62.53 64.56 55.01 46.06 58.32 Qwen2-VL 50.50 48.88 42.38 46.46 51.71 47.99 FAK-Owl 67.54 60.85 67.99 69.84 70.72 67.39 Ours-LVLM 97.11 89.90 92.26 95.97 86.70 92.39
Table 3: Comparison of our scheme with existing LVLM-based methods. Clipping (Khan & Dang-Nguyen, 2024) and PandaGPT (Su et al., 2023) are fine-tuned on the FF++ dataset, FAK-Owl (Liu et al., 2024b) is trained based on original image-text pairs, and Qwen2-VL (Wang et al., 2024a) is evaluated with its original pre-trained weights. The AUC scores for Ours-LVLM are derived from the textual output of the LLM, and the AUC scores for Ours-KFD are derived from the KFD output. The best and second-best results are highlighted in bold and underlined, respectively.
Detection Performance of LVLM. We benchmark our framework against state-of-the-art LVLM-based classification methods (Khan & Dang-Nguyen, 2024; Lin et al., 2025; Fu et al., 2025) and VQA methods (Su et al., 2023; Wang et al., 2024a; Liu et al., 2024b). In these evaluations, both the image and the corresponding query are provided as inputs to the LVLM, and the model is required to determine the authenticity of the image (real or fake). For classification task, our Knowledge-guided Forgery Detector (KFD) demonstrates significantly superior detection performance compared to existing LVLM-based classification methods. For VQA models such as PandaGPT, Qwen2-VL, and FAK-Owl, we utilize the LLM’s output (“Yes” or “No”) to classify authenticity and subsequently compute the AUC values for evaluation. As summarized in Table 3, our method consistently outperforms existing LVLM-based VQA methods, achieving better performance on the FF++, CDF2, DFD, CDF1, and DFDCP datasets.
Figure 5: Forgery Localization and LLM Responses under Intra-dataset and Cross-dataset Evaluations. Each example includes the original image (top-left), the corresponding forgery segmentation map (top-right), and the textual detection result generated by the LLM (bottom).
Dialogues Visualization. Unlike prior detection methods, our approach not only supports deepfake detection but also facilitates multi-turn dialogues, enabling users to further inquire about the image content. Figure 5 presents some example dialogues under intra-dataset and cross-dataset evaluations. It shows that our proposed scheme accurately identifies forged regions within the images, while the LLM provides precise and contextually relevant judgments. Additional multi-turn dialogue examples are provided in the supplementary material.
@@ 4.4 Analysis
Number of Training Images. Obtaining large-scale face images in real-world scenarios is often infeasible. Hence, we assess our model’s performance with varying numbers of training images. Specifically, we randomly sample 50, 100, 200, and 500 real images from FF++ to create corresponding fake image-text pairs for training. The training steps are fixed as 500. The video-level AUC is calculated based on the responses of LLM. As shown in Table 4, our approach achieves state-of-the-art performance using only 500 training images. Although reducing the training set size leads to a slight degradation in accuracy, our method still maintains competitive results with as few as 100 training images. This underscores the robustness of our framework, particularly in scenarios where data are scarce.
Effect of Prompt Tuning. The prompt tuning process is designed to convert forgery detection knowledge into the input of LLM to facilitate accurate detection. This process involves the Forgery Prompt Learner (FPL), the LLM, and the LoRA strategy. To evaluate the effectiveness of this component, we conduct ablation experiments on FF++, CDF2, and DFDC datasets. For each configuration, the AUC is calculated based on the LLM’s output in determining real versus fake images. As shown in Table 5, models equipped with the Forgery Prompt Learner demonstrate higher AUC values, indicating improved effectiveness for deepfake detection tasks. Furthermore, the integration of LoRA further enhances performance, achieving superior results across multiple datasets compared to configurations without LoRA.
Number Test set AUC CDF1 CDF2 DFD DFDC DFDCP 50 77.78 82.64 82.25 70.49 84.70 100 82.59 81.74 86.34 72.72 85.06 200 84.53 80.68 88.14 74.12 83.48 500 84.17 83.31 88.17 74.23 84.65
@@
Training Set Test Set AUC (%) Dataset #Real FF++ CDF2 DFDCP DFDC Avg FF++ 720 99.54 94.32 92.39 77.03 90.82 CDF2 622 97.95 95.79 86.36 70.27 87.59 DFDCP 737 99.15 85.17 93.74 69.64 86.93
@@
LLM Architecture Test set AUC (%) FF++ CDF1 DFD DFDC Avg Llama-3.2-1B 96.49 95.64 82.74 65.66 85.13 Llama-3.2-3B 97.08 95.30 82.80 66.12 85.32 Vicuna-7B 97.11 95.97 84.26 67.65 86.25
@@
Generalizability to Training Datasets. The generalizability of deepfake detection is closely related to the diversity of the training data used. To verify the effectiveness of our approach across different datasets, we train the model on various training sets and conducted cross-dataset evaluation on FF++, CDF2, DFDCP, and DFDC datasets. We calculate detection AUC values based on the LLM’s responses, as shown in Table 6. The results indicate that our method exhibits strong robustness across diverse datasets and demonstrates its ability to generalize across different types of forgeries using varied real data.
Effects of different LLM architectures. The deepfake detection performance is also related to the specific LLM architecture used, as different LLMs exhibit unique characteristics. To examine detection performance across various LLM architectures, we evaluate three models: Llama-3.2-1B, Llama-3.2-3B, and Vicuna-7B. The evaluation is conducted on the FF++, CDF1, DFD, and DFDC datasets. As shown in Table 7, we observe that larger-scale LLM architectures consistently yield superior detection performance. This trend suggests that models with increased parameter capacity are more adept at capturing fine-grained forgery artifacts.
Ablation Study on the Reference-based Optimization. To validate the effectiveness of the reference based optimization process, we train the model under two configurations: with and without Reference-based Optimization Process (ROP). We then evaluate the generalization capability of our approach across various datasets. Table 8 summarizes the generalization performance on the CDF1, CDF2, DFDC, and DFDCP datasets. The results demonstrate that even without similarity optimization, the proposed framework outperforms existing methods. Furthermore, introducing the similarity optimization process yields additional performance gains, underscoring its effectiveness in enhancing the generalization capability of deepfake detection models.
Methods Test set AUC (%)
CDF1 CDF2 DFDC DFDCP Avg
SBI 93.44 93.82 74.47 90.95 88.17
w/o ROP 96.18 92.77 79.31 91.73 90.00
w/ ROP 97.62 94.71 79.12 91.81 90.82
Table 8: Ablation study on the Reference-based Optimization Process. The highest AUC scores among all methods are highlighted in bold.
@@ 5 Conclusion
In this work, we introduce a novel deepfake detection framework that leverages LVLMs to enhance generalization and explainability. By integrating a Knowledge-guided Forgery Detector, we effectively align image features with textual descriptions of pristine and deepfake images to facilitate forgery classification and localization. Furthermore, we incorporate a Forgery Prompt Learner capable of transforming fine-grained forgery features into inputs for the LLM, enabling accurate forgery detection responses. Extensive evaluations across multiple benchmarks, including FF++, CDF1, CDF2, DFD, DFDCP, and DFDC, demonstrate that our scheme outperforms existing methods in generalization performance. Notably, our framework also supports multi-turn dialogue, providing interactive and explainable detection results. These findings underscore the potential of LVLM-based approaches in advancing deepfake detection methodologies.
Acknowledgements
This work is supported in part by the National Key R&D Program of China under Grant number 2022YFB3103100, in part by the National Natural Science Foundation of China under grant numbers, U23B2023 and 62472199, Guangdong Key Laboratory of Data Security and Privacy Preserving under Grant 2023B1212060036, in part by Engineering Research Center of Trustworthy AI, Ministry of Education, in part by the Ministry of Education, Singapore, under its Academic Research Fund (AcRF) Tier 2 Award No. MOET2EP50220-0003.
Impact Statement
This research aims to improve generalizability and explainability for deepfake detection. However, the reliance on large-scale datasets for training raises potential privacy concerns, and biases in the training data could lead to unequal performance across different demographic groups. Nevertheless, developing algorithms capable of detecting deepfake data is essential, and we advocate for robust security measures and legal frameworks to govern the use of such technology.
Appendix A Appendix
A.1 Implementation Details
Knowledge-guided Forgery Detector (KFD): To incorporate external knowledge for precise deepfake detection, we leverage the image and text encoders of the ImageBind-huge model to extract image content features and textual description features. For the visual features, we select layers 16, 24, and 32 from the image encoder to extract corresponding features. Since the dimensionality of the text features is fixed at 4096, while the dimensionality of the image features varies across layers, we employ linear layers to project the image content features to match the dimensionality of the text features. By calculating the correlation between the text features and each set of image content features, we generate three consistency maps. These maps are subsequently passed to the Forgery Locator and Classifier for forgery localization and classification. During training, the batch size is set to 16.
Real and fake descriptions: Textual descriptions are generated via GPT-4, validated by human annotators. Examples include “Inconsistent head poses” or “Mismatched skin texture”. Detailed examples are provided as follows:
deepfake_descriptions = ["unnatural blending edges", "inconsistent head poses", "noticeable blending artifacts", "mismatched skin texture", "unnatural reflections", "unnatural looking hairlines", "misaligned facial features", "digital artifacts visible", "strange light flares around the lips", "jawline appears overly smooth", "distorted shadows", "over-sharpened or exaggerated facial features", "distorted edges around the face", "fake"]
real_face_descriptions = ["natural light and shadow transitions", "natural eye reflections", "natural appearance", "natural skin tones and color variations", "realistic reflections in their eyes", "naturally textured skin", "fine hair near the eyebrows and forehead", "realistic texture", "natural sparkle in their eyes", "natural highlights on the bridge of the nose", "natural skin texture", "natural skin folds around the eyes", "natural light sources", "smooth yet textured skin on the forehead", "skin texture is detailed and realistic", "real"]
About LLM: Due to computational constraints, we employ Vicuna-7B as the LLM to process visual, forgery, and question prompt embeddings. For LoRA adaptation, we add low-rank matrices with a rank of 32 to the $q\_proj$, $k\_proj$, $v\_proj$, and $o\_proj$ modules. The $lora\_alpha$ parameter is set to 32, and $lora\_dropout$ is set to 0.1. During LLM training, the batch size is set to 1.
Video-level AUC Calculation for LLM: The video-level AUC is computed by aggregating frame-level binary outputs. For each video, we sample 32 frames uniformly and calculate the ratio of “yes” responses (indicating “fake”). This ratio indicates the video’s probability score to be a fake. To extract “yes” response from model output, we implemente a deterministic rule-based parsing strategy for extracting binary labels (“yes”/“no”) from model outputs. If the output contains “yes” or “is deepfake”, the frame is labeled fake. If the response contains “no” or “not deepfake”, the frame is labeled real. If none of the keywords are detected, the response is default labeled real.
A.2 Robustness to Perturbations
Given that deepfake content is frequently shared on social media, it is often subject to various perturbations such as noise, compression, and image enhancement. To assess robustness under these conditions, we apply a range of perturbations to the FF++ dataset and then evaluate detection AUC and text description accuracy. As shown in Table 9, we follow the settings from (Nguyen et al., 2024) to benchmark our approach against existing methods. The results demonstrate that our method achieves better robustness compared to prior approach (Shiohara & Yamasaki, 2022), maintaining high AUC values even under challenging conditions.
Methods Saturation Constrast Block JPEG Noise Blur
SBI 99.13 98.37 99.15 77.88 58.10 67.12
Ours 99.20 98.12 99.54 87.96 62.69 81.37
Table 9: Robustness evaluation on FF++. The highest AUC values among all methods are highlighted in bold.
A.3 Inference Time
In this section, we compare the inference time of our approach with existing approaches using a single NVIDIA RTX 4090 GPU. The inference time, measured in seconds per frame, is reported alongside the corresponding AUC scores for reference. As shown in Table 10, the variant employing only KFD exhibits a slightly longer inference time than CADDM, yet achieves substantially higher precision. Similarly, our LVLM incurs a greater inference time than FAK-Owl but also attains significantly superior precision. Moreover, unlike FAK-Owl which is constrained to binary (Yes/No) responses, our approach supports multi-turn dialogues providing improved interpretability and enhanced generalization. Overall, despite a higher inference time, our approach provides more precise detection and comprehensive interpretability.
Table 10: Inference time and AUC on CDF2 dataset.
Inference time per frame(s) AUC
CADDM 0.026 85.68
Ours-KFD 0.059 97.62
FAK-Owl 0.642 69.84
Ours-LVLM 1.211 95.97
A.4 Failure Cases and Limitations
Our approach faces limitations in the training strategy. The alternating training strategy for multi-turn dialogue introduces domain gaps: general-purpose VQA datasets prioritize object-centric reasoning, whereas fine-grained forgery detection requires localized artifact analysis. This misalignment occasionally results in a decrease in forgery detection performance (see Table 3). The Failure cases are shown in Figure 6. To address these limitations, we will construct domain-specific forgery QA datasets with spatially grounded annotations in future works.
Figure 6: Failure cases of our method. While the model precisely locate forgery regions, the LLM fails to generate accurate textual descriptions corresponding to the identified content.
A.5 Conversations under various datasets
In this section, we present additional dialogue results. We apply our method across five subsets of the FF++ dataset and obtain the corresponding responses, as shown in Figure 7. It can be observed that the responses accurately determine authenticity. Furthermore, as illustrated in Figure 8, our proposed approach continues to effectively identify forgeries in cross-dataset conversations.
Figure 7: Forgery localization results and LLM responses on various forgery types from the FF++ dataset.
Figure 8: Forgery localization results and LLM responses on various forgery types from the CDF2, DFDCP, DFDC, and DFD datasets.
A.6 Multi-turn Conversations
Unlike traditional deepfake detection algorithms, our proposed framework not only achieves accurate detection but also supports multi-turn dialogue capabilities. For instance, when a user queries, “Is this a deepfake image?”, the model provides precise responses, identifying whether the image is manipulated and offering relevant details. Furthermore, the framework enables users to inquire about additional information, such as forgery scores or specific aspects of the image content. As illustrated in Figures 926, we present examples from the FF++, CDF1, CDF2, and DFDC datasets. These results demonstrate the model’s ability to generate accurate and contextually relevant responses to user queries, underscoring its potential for real-world applications.
Figure 9: Conversations with real images from the FaceForensics++ dataset.
Figure 10: Conversations with real images from the FaceForensics++ dataset.
Figure 11: Conversations with real images from the FaceForensics++ dataset.
Figure 12: Conversations with real images from the FaceForensics++ dataset.
Figure 13: Conversations with DeepFake images from the FaceForensics++ dataset.
Figure 14: Conversations with DeepFake images from the FaceForensics++ dataset.
Figure 15: Conversations with Face2Face images from the FaceForensics++ dataset.
Figure 16: Conversations with Face2Face images from the FaceForensics++ dataset.
Figure 17: Conversations with FaceSwap images from the FaceForensics++ dataset.
Figure 18: Conversations with FaceSwap images from the FaceForensics++ dataset.
Figure 19: Conversations with NeuralTexture images from the FaceForensics++ dataset.
Figure 20: Conversations with NeuralTexture images from the FaceForensics++ dataset.
Figure 21: Conversations with fake images from the CDF1 dataset.
Figure 22: Conversations with fake images from the CDF1 dataset.
Figure 23: Conversations with fake images from the CDF2 dataset.
Figure 24: Conversations with fake images from the CDF2 dataset.
Figure 25: Conversations with fake images from the DFDC dataset.
Figure 26: Conversations with fake images from the DFDC dataset.